OpenStudy (anonymous):

Help Use the following function. log4(x + 3) = 2 The second step would be: Select one: a. Take the square root of each side. b. Take the log of each side. c. Divide each side by 16. d. To simplify the equation using inverses to isolate for x.

OpenStudy (zarkon):

you might want to clarify if your equation is \[\log4(x+3)=2\]or \[\log_{4}(x+3)=2\]

OpenStudy (anonymous):

the 2nd one sorry

OpenStudy (michele_laino):

hint: using the definition of logarithm, we can write: \[\huge 2 = {\log _4}16\]

OpenStudy (campbell_st):

ok... so if you raise each term as a power of the base then you will be able to solve it if \[\log_{a}(b) = n\] then \[a^{\log_{a}(b)} = a^n\] which can be simplified to \[b = a^n\] you need to apply this rule for logs to your porblem where a = 4, b = x + 3 and n = 2 hope it helps

OpenStudy (anonymous):

okay so once we write that @Michele_Laino what do we do next?

OpenStudy (michele_laino):

using my identity above, we can rewrite your equaton as follows: \[\huge {\log _4}\left( {x + 3} \right) = {\log _4}16\]

OpenStudy (michele_laino):

then we can equate the numbers of both logarithm, like below: \[\huge x + 3 = 16\] please solve for \(x\)

OpenStudy (anonymous):

13

OpenStudy (anonymous):

i still dont understand for the answer .. what was the 2nd step?

OpenStudy (michele_laino):

that's right! So, what is the right option?

OpenStudy (anonymous):

B?

OpenStudy (michele_laino):

as we can see, when I wrote my identity, I took the logarithm of both sides, so you are right, it is option B

OpenStudy (anonymous):

thanks can you help me again?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Use the following equation: log2x + log2(x – 7) = 3 The second step would be to: Select one: a. Divide both sides by 3 to simplify. b. Re-write the equation using the definition of logarithms. c. Add (x – 7) to both sides to simplify. d. Either A or C

OpenStudy (michele_laino):

If I use the definition of logarithm, I can write this: \[\huge 3 = {\log _2}8\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so the 2nd step would be Re-write the equation using the definition of logarithms.

OpenStudy (michele_laino):

after that it is simple to solve the logarithm equation. So what is the right option?

OpenStudy (michele_laino):

that's right! It is option B

OpenStudy (anonymous):

Hey i need help in another one i think i got the answer to it but i need to make sure ill tag you in it ??

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

hold on!